A stream of electrons, each of energy 𝑬 = πŸ‘π’†π‘½ is incident on a potential of height π‘½πŸŽ = πŸ’π’†π‘½ . The width of the barrier is 20Γ…. Calculate the percentage transmission of the beam through this barrier.?


1
Expert's answer
2021-03-01T12:38:51-0500

"E=3eV"

"V_{o}=4eV"

"L=20A^o=20\times10^{-10}m"

"\alpha L=\sqrt{\dfrac{2m(V_o-E)}{\hbar}}L\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space m=9.1\times10^{-31}kg"

"\alpha L=10.23"


Transmittance, "T=16\times\dfrac{3}{4}\times(1-\dfrac{3}{4})\times e^{-2\alpha L}"

"T=3.90\times10^{-9}"